The time taken to assemble a car in certain plant is a random variable having a normal distribution of 20 hours and a standard deviation of 2 hours. What is the probability the car can be assembled at this plant in a period of time

a) less than 19.5 hours?

b) between 20 and 22 hours?


1
Expert's answer
2022-02-02T12:03:17-0500

Let X denote the time taken to assemble a car in a certain plant. Given "\mu =20, \sigma=2".

Let "Z = \dfrac{X - 20}{2}".

a)

"\begin{aligned}
P(X<19.5) &= P\left(\dfrac{X - 20}{2} < \dfrac{19.5 -20}{2}\right)\\
&= P(Z < -0.25)\\
&= 1 - P(Z<0.25)\\
&= 1 - 0.5987 ~~(\text{Using Standard Normal Table})\\
&=0.4013
\end{aligned}"

b)

"\begin{aligned}
P(20<X<22) &= P\left(\dfrac{20-20}{2}<\dfrac{X - 20}{2} < \dfrac{22 -20}{2}\right)\\
&= P(0< Z < 1)\\ &= P(Z<1)-P(Z<0)\\
&=0.8413-0.5\\
&= 0.3413 ~~(\text{Using Standard Normal Table})\\
\end{aligned}"